# Towards Efficient Training of Graph Neural Networks: A Multiscale Approach

### v3

@@ Abstract
Graph Neural Networks (GNNs) have become powerful tools for learning from graph-structured data, finding applications across diverse domains. However, as graph sizes and connectivity increase, standard GNN training methods face significant computational and memory challenges, limiting their scalability and efficiency. In this paper, we present a novel framework for efficient multiscale training of GNNs. Our approach leverages hierarchical graph representations and subgraphs, enabling the integration of information across multiple scales and resolutions. By utilizing coarser graph abstractions and subgraphs, each with fewer nodes and edges, we significantly reduce computational overhead during training. Building on this framework, we propose a suite of scalable training strategies, including coarse-to-fine learning, subgraph-to-full-graph transfer, and multiscale gradient computation. We also provide some theoretical analysis of our methods and demonstrate their effectiveness across various datasets and learning tasks. Our results show that multiscale training can substantially accelerate GNN training for large scale problems while maintaining, or even improving, predictive performance.
@@ 1 Introduction
Large-scale graph networks arise in diverse fields such as social network analysis Kipf & Welling (2016); Defferrard et al. (2016), recommendation systems Tang et al. (2018), and bioinformatics Jumper et al. (2021), where relationships between entities are crucial to understanding system behavior Abadal et al. (2021). These problems involve processing large graph datasets, necessitating scalable solutions. Some examples of those problems include identifying community structures in social networks Hamilton et al. (2017), optimizing traffic flow in smart cities Geng et al. (2019), or studying protein-protein interactions in biology Xu et al. (2024). In all these problems, efficient processing of graph data is required for practical reasons.
In recent years, many Graph Neural Networks (GNNs) algorithms have been proposed to solve large graph problems, such as GraphSage Hamilton et al. (2017) and FastGCN Chen et al. (2018). However, despite their success, when the involved graph is of a large scale, with many nodes and edges, it takes a long time and many resources to train the network Galmés et al. (2021).
In this paper, we study and develop multiscale training methodologies, to improve the training efficiency of GNNs. To achieve that, we focus on the standard approach for implementing GNNs, that is, the Message-Passing (MP) scheme Gilmer et al. (2017). Mathematically, this approach is equivalent to the multiplication of the graph adjacency matrix ${\bf A}$ with node feature maps ${\bf X}^{(l)}$, which is then multiplied by a learned channel-mixing weight matrix ${\bf W}^{(l)}$, followed by an application of an element-wise nonlinear activation function $\sigma$, as follows:
@@
Assuming that the adjacency matrix is sparse with an average sparsity of $k$ neighbors per node, and a channel-mixing matrix of shape $c\times c$, we now focus on the the total multiplications in the computation of Equation (1), to which we refer to as ’work’. In this case, the work required in a single MP layer is of order ${\cal O}(n\cdot k\cdot c^{2})$, where $n$ is the number of the nodes in the graph, and $c$ is the number of channels. For large or dense graphs, i.e., $n$ or $k$ are large, the MP in Equation (1) requires a high computational cost. Any algorithm that works with the entire graph at once is therefore bound to face the cost of this matrix-vector product.
@@
A possible solution to this high computational cost involves the utilization of a multiscale framework, which requires the generation of surrogate problems to be solved (Kolaczyk & Nowak, 2003). If the solution of the surrogate problem approximates the solution of the fine (original) scale problem, then it can be used as an initial starting point, thus reducing the work required to solve the optimization problem. In the context of graphs, such a multiscale representation of the original problem can be achieved using graph coarsening. Graph coarsening algorithms are techniques that reduce the size of graphs while preserving their essential properties (Cai et al., 2021). These methods aim to create smaller, more computationally manageable graphs by choosing a subset of the nodes or the edges, based on certain policies, or by using a subgraph of the original graph, and may be implemented in various ways.
@@
We demonstrate the generation of a coarse graph using random pooling and Topk pooling (which is based on the nodes with high degrees in the graph) in Figure 1. As for choosing a subgraph, this can be implemented using node-based selection methods, such as ego-networks (Gupta et al., 2014) or node marking and deletion (Frasca et al., 2022), which offer straightforward ways to extract subgraphs based on specific nodes of interest. We show an example of subgraph coarsening in Figure 2.
Our approach.
@@
While coarsening and subgraph methods have been in use for various purposes, from improved architectures Gao & Ji (2019), to enhanced expressiveness Bevilacqua et al. (2022; 2024), to the best of our knowledge, they were not considered in the context of efficient training of GNNs. Therefore, in this work, we explore three novel training mechanisms to design surrogate processes that approximate the original training problem with a fraction of the computational cost, using a multiscale approach. All the methods attempt to use the full graph with $n$ nodes as little as possible, and focus on training the network using coarse-scale graphs, thus reducing computational cost. The three methods, detailed in Section 4, are as follows:
@@
  • Coarse-to-Fine. A coarse-to-fine algorithm approximates the optimization problem on smaller, pooled graphs. If the coarse-scale solution is similar to the fine-scale one, it can serve as an effective initialization for the fine-scale optimization process. Since gradient-based methods converge faster when initialized near the solution (Arora et al., 2018; Boyd & Vandenberghe, 2004), this approach reduces the overall number of iterations required.
@@
  • Sub-to-Full. This approach leverages a sequence of subgraphs that progressively grow to the original fine-scale graph. Like coarsening, each subgraph is significantly smaller than the original graph, enabling faster optimization process. Subgraphs can be generated based on node distances (if available) or ego-network methods Gupta et al. (2014).
@@
  • Multiscale Gradients Computation. Building on the previous methods and inspired by Multiscale Monte Carlo Giles (2001), which is widely used for solving stochastic differential equations, we introduce a mechanism to approximate fine-scale gradients at a fraction of the computational cost, achieved by combining gradients across scales to efficiently approximate the fine-scale gradient.
Contributions.
The main contribution of this paper is the design of novel graph multiscale optimization algorithms to obtain computationally efficient GNN training methods that retain the performance of standard gradient-based training of GNNs. To demonstrate their effectiveness, we benchmark the obtained performance and reduce computational costs on multiple graph learning tasks, from transductive node classification to inductive tasks like shape segmentation.
@@ 2 Related Work
We discuss related topics to our paper, from hierarchical graph pooling to multiscale methods and their learning.
@@
Graph pooling can be performed in various ways, as reviewed in Liu et al. (2022). Popular pooling techniques include DiffPool (Ying et al., 2018) MaxCutPool (Abate & Bianchi, 2024), graph wavelet compression Eliasof et al. (2023), as well as Independent-Set Pooling (Stanovic et al., 2024), all aiming to produce a coarsened graph that bears similar properties to the original graph. Wang et al. (2020) and Wang et al. (2024) introduce the concept of subgraph pooling to GNNs. Bianchi & Lachi (2023) study the expressive power of pooling in GNNs. Here, we use graph pooling mechanisms, as illustrated in Figure 1Figure 2, as core components for efficient training.
@@
The idea of Coarse-to-Fine for the solution of optimization problems has been proposed in Brandt (1984), while a more detailed analysis of the idea in the context of optimization problems was introduced in Nash (2000). Multiscale approaches are also useful in the context of Partial Differential Equations (Guo & Li, 2024; Li et al., 2024b), as well as for various applications including flood modeling (Bentivoglio et al., 2024). Lastly, Monte-Carlo methods (Hastings, 1970) have also been adapted to a multiscale framework (Giles, 2001), which can be used for various applications such as electrolytes and medical simulations (Liang et al., 2015; Klapproth et al., 2021). In this paper, we draw inspiration from multiscale methods to introduce a novel and efficient GNN training framework.
@@
In the context of graphs, the recent work in Shi et al. (2022) has shown that a matching problem can be solved using a coarse-to-fine strategy, and Cai et al. (2021) proposed different coarsening strategies. Other innovations include the recent work of Li et al. (2024a), whose goal is to mitigate the lack of labels in semi-supervised and unsupervised settings. We note that it is important to distinguish between algorithms that use a multiscale structure as a part of the network, such as Unet (Ronneberger et al., 2015) and Graph-Unet (Gao & Ji, 2019), and multiscale training that utilizes multiscale for training. In particular, the latter, which is the focus of this paper, can be potentially applied to any GNN architecture.
@@ 3 Preliminaries and Notations
We now provide related graph-learning notations and preliminaries that we will use throughout this paper. Let us consider a graph defined by ${\cal G=(\cal V,\cal E)}$ where $\cal V$ is a set of $n$ nodes and $\cal E$ is a set of $m$ edges. Alternatively, a graph can be represented by its adjacency matrix ${\bf A}$, where the $(i,j)$-th entry is 1 if an edge between nodes $i$ and $j$ exists, and 0 otherwise. We assume that each node $i$ is associated with a feature vector with $c$ channels, ${\bf x}_{i}\in\mathbb{R}^{c}$, and ${\bf X}=[{\bf x}_{0}\ldots,{\bf x}_{n-1}]\in\mathbb{R}^{n\times c}$ is the node features matrix. We denote by ${\bf y}_{i}\in\mathbb{R}^{d}$ the $i$-th node label, and by ${\bf Y}=[{\bf y}_{0},\ldots,{\bf y}_{n-1}]\in\mathbb{R}^{n\times d}$ the label matrix.
@@
For instance, in the well-known Graph Convolution Network (GCN) (Kipf & Welling, 2016), each layer is defined using Equation (1), such that ${\bf X}^{(l+1)}$ is the output of the $l$-th layer and serves as the input for the $(l+1)$-th layer. For a network with $L$ layers, we denote by ${\boldsymbol{\theta}}=\{{\bf W}^{(1)},\ldots,{\bf W}^{(L)}\}$ the collection of all learnable weight matrices within the network.
@@
Lastly, we note that, as shown in Section 6, our approach is general and applicable to various GNN architectures.
@@ 4.1 Coarse-to-Fine Training
One seemingly simple idea is to solve the $r$-th surrogate problem in Equation (4) and use the obtained weights to solve the problem on the $(r-1)$-th graph, that is ${\cal G}_{{r-1}}$, which is larger. This approach can be beneficial only if $\widehat{\boldsymbol{\theta}}_{r+1}\approx\widehat{\boldsymbol{\theta}}_{r}$, that is, the obtained optimal weights are similar under different scales, and if the total amount of work on the coarse scale is smaller than on the fine one, such that it facilitates computational cost reduction. The multiscale framework is summarized in Algorithm 1.
@@
While Algorithm 1 is relatively simple, it requires obtaining the $r$-th surrogate problem by coarsening the graph, as well as a stopping criterion for the $r$-th optimization problem. Generally, in our coarse-to-fine method, we use the following graph coarsening scheme:
@@
Choosing the non-zero entries of the matrix ${\bf P}_{r}$, which are the nodes that remain in the graph after pooling, can be implemented in various ways. For example, ${\bf P}_{r}$ can be obtained by sampling random nodes of the fine graph $\mathcal{G}$, as illustrated in Figure 1 (left), or with some desired pooling logic, such as choosing nodes of the highest degree in the graph, which is a method often described as $Topk$ pooling (Huang et al., 2015) and we present an illustration of it in Figure 1 (right). The power $p$ in Equation (5) controls the connectivity of the graph, where $p=1$ is the original graph adjacency matrix, and using a value $p>1$ will enhance the connectivity of the graph prior to the coarsening step.
In Algorithm 1, we apply GNN layers such as in Equation (1), using the coarsened feature matrix ${\bf X}_{r}$ and its corresponding adjacency matrix ${\bf A}_{r}$, that is based on a desired pooling ratio. We note that there is no re-initialization of the weight matrix ${\bf W}$ when transferring from coarse grids to finer ones, allowing a continuous learning process along the steps of the algorithm. As we show later in our experiments in Section 6, our coarse-to-fine training approach, summarized in Algorithm 1, is not restricted to a specific GNN architecture or a pooling technique.
@@ 4.2 Sub-to-Full Training
We now describe an additional method for efficient training on GNNs, that relies on utilizing subgraphs, called Sub-to-Full. At its core, the method is similar to Coarse-to-Fine, as summarized in Algorithm 1, using the same coarsening in Equation (5). However, its main difference is that we now choose the indices of ${\bf P}_{r}$ such that the nodes ${\bf X}_{r}$ form a subgraph of the original graph $\cal G$.
A subgraph can be defined in a few possible manners. One is by choosing a root node $j$, and generating a subgraph centered around that node. If the graph $\cal{G}$ has some geometric features, meaning each input node corresponds to a point in the Euclidean space $\mathbb{R}^{d}$, one might consider constructing a subgraph with $n_{r}$ nodes to be the $n_{r}$ nodes that are closest to the chosen center node in terms of Euclidean distance norm. Otherwise, we can consider a subgraph created using an ego-network Gupta et al. (2014) of $k-hop$ steps, such that we consider all neighbors that are $k$ steps away from the central node. An illustration of this method is presented in Figure 2. The number of hops, $k$, is a hyperparameter of the algorithm process, and we increase it as we get from coarse levels to fine ones within the series of graph $\{\mathcal{G}_{r}\}_{r=1}^{R}$.

### v5

@@ Abstract
Graph Neural Networks (GNNs) have become powerful tools for learning from graph-structured data, finding applications across diverse domains. However, as graph sizes and connectivity increase, standard GNN training methods face significant computational and memory challenges, limiting their scalability and efficiency. In this paper, we present a novel framework for efficient multiscale training of GNNs. Our approach leverages hierarchical graph representations and subgraphs, enabling the integration of information across multiple scales and resolutions. By utilizing coarser graph abstractions and subgraphs, each with fewer nodes and edges, we significantly reduce computational overhead during training. Building on this framework, we propose a suite of scalable training strategies, including coarse-to-fine learning, subgraph-to-full-graph transfer, and multiscale gradient computation. We also provide some theoretical analysis of our methods and demonstrate their effectiveness across various datasets and learning tasks. Our results show that multiscale training can substantially accelerate GNN training for large-scale problems while maintaining, or even improving, predictive performance.
@@ 1 Introduction
Large-scale graph networks arise in diverse fields such as social network analysis (gcn; defferrard_2016), recommendation systems (recommendation_systems), and bioinformatics (jumper2021highly), where relationships between entities are crucial to understanding system behavior (abadal2021computing). These problems involve processing large graph datasets, necessitating scalable solutions. Some examples of those problems include identifying community structures in social networks (hamilton2017inductive), optimizing traffic flow in smart cities (ride_hailing_gcn), or studying protein-protein interactions in biology (protein_protein). In all these problems, efficient processing of graph data is required for practical reasons.
In recent years, many Graph Neural Networks (GNNs) algorithms have been proposed to solve large graph problems, such as GraphSage (hamilton2017inductive) and FastGCN (chen2018fastgcn). However, despite their success, when the involved graph is of a large scale, with many nodes and edges, it takes a long time and many resources to train the network (upscaling_gnns).
In this paper, we study and develop multiscale training methodologies, to improve the training efficiency of GNNs. To achieve that, we focus on the standard approach for implementing GNNs, that is, the Message-Passing (MP) scheme (gilmer2017neural). Mathematically, this approach is equivalent to the multiplication of the graph adjacency matrix $\bf A$} with node feature maps ${\bf X}^{(l)}$, which is then multiplied by a learned channel-mixing weight matrix ${\bf W}^{(l)}$, followed by an application of an element-wise nonlinear activation function $\sigma$, as follows:
@@
Assuming that the adjacency matrix is sparse with an average sparsity of $k$ neighbors per node, and a channel-mixing matrix of shape $c\times c$, we now focus on the total multiplications in the computation of Equation (1), to which we refer to as ’work’. In this case, the work required in a single MP layer is of order ${\cal O}(n\cdot k\cdot c^{2})$, where $n$ is the number of nodes in the graph, and $c$ is the number of channels. For large or dense graphs, i.e., $n$ or $k$ are large, the MP in Equation (1) requires a high computational cost. Any algorithm that works with the entire graph at once is therefore bound to face the cost of this matrix-vector product.
@@
A possible solution to this high computational cost involves the utilization of a multiscale framework, which requires the generation of surrogate problems to be solved (Kolaczyk2003). If the solution of the surrogate problem approximates the solution of the fine (original) scale problem, then it can be used as an initial starting point, thus reducing the work required to solve the optimization problem. In the context of graphs, such a multiscale representation of the original problem can be achieved using graph coarsening. Graph coarsening algorithms are techniques that reduce the size of graphs while preserving their essential properties (cai2021graph). These methods aim to create smaller, more computationally manageable graphs by choosing a subset of the nodes or the edges, based on certain policies, or by using a subgraph of the original graph, and may be implemented in various ways.
@@
We demonstrate the generation of a coarse graph using random pooling and Topk pooling (which is based on the nodes with high degrees in the graph) in Figure˜1. As for choosing a subgraph, this can be implemented using node-based selection methods, such as ego-networks (ego_networks) or node marking and deletion (subgarph_gnns_fresca), which offer straightforward ways to extract subgraphs based on specific nodes of interest. We show an example of subgraph coarsening in Figure˜2.
Contribution.
@@
In this paper, we propose novel graph multiscale optimization algorithms to obtain computationally efficient GNN training methods that retain the performance of standard gradient-based training of GNNs. Our approach is based on a multiscale framework that utilizes various graph coarsening techniques. While coarsening and subgraph methods have been in use for various purposes, from improved architectures (graph_u_net), to enhanced expressiveness (bevilacqua2022equivariant; bevilacqua2024efficient), to the best of our knowledge, they were not considered in the context of efficient training of GNNs. Therefore, in this work, we explore three novel training mechanisms to design surrogate processes that approximate the original training problem with a fraction of the computational cost, using a multiscale approach. All the methods attempt to use the full graph with $n$ nodes as little as possible, and focus on training the network using coarse-scale graphs, thus reducing computational cost. The three methods, detailed in Section˜4, are as follows:
@@
  • Coarse-to-Fine. A coarse-to-fine algorithm approximates the optimization problem on smaller, pooled graphs. If the coarse-scale solution is similar to the fine-scale one, it can serve as an effective initialization for the fine-scale optimization process. Since gradient-based methods converge faster when initialized near the solution (arora2018convergence; boydBook), this approach reduces the overall number of iterations required.
@@
  • Sub-to-Full. This approach leverages a sequence of subgraphs that progressively grow to the original fine-scale graph. Like coarsening, each subgraph is significantly smaller than the original graph, enabling faster optimization process. Subgraphs can be generated based on node distances (if available) or ego-network methods (ego_networks).
@@
  • Multiscale Gradients Computation. Building on the previous methods and inspired by Multiscale Monte Carlo (mlcm), which is widely used for solving stochastic differential equations, we introduce a mechanism to approximate fine-scale gradients at a fraction of the computational cost, achieved by combining gradients across scales to efficiently approximate the fine-scale gradient.
To demonstrate their effectiveness, we benchmark the obtained performance and reduce computational costs on multiple graph learning tasks, from transductive node classification to inductive tasks. Our code is available here: https://github.com/eshedgal1/GraphMultiscale.
@@ 2 Related Work
We discuss related topics to our paper, from hierarchical graph pooling to multiscale methods.
@@
Graph pooling can be performed in various ways, as reviewed in graph_pooling. Popular pooling techniques include DiffPool (graph_rep_dif_pooling) MaxCutPool (max_cut_pul), graph wavelet compression (eliasof2023haar), as well as Independent-Set Pooling (independent_set_pooling), all aiming to produce a coarsened graph that bears similar properties to the original graph. subgraph_cropping and subgraph_pooling introduce the concept of subgraph pooling to GNNs. power_of_pooling study the expressive power of pooling in GNNs. Here, we use graph pooling mechanisms, as illustrated in Figure˜1Figure˜2, as core components for efficient training.
@@
The idea of Coarse-to-Fine for the solution of optimization problems has been proposed in brandt, while a more detailed analysis of the idea in the context of optimization problems was introduced in NashMGOPT. Multiscale approaches are also useful in the context of Partial Differential Equations (multigrid_pde; harnessing_scale_and_physics), as well as for various applications, including flood modeling (flood_modeling). Lastly, Monte-Carlo methods (monte_carlo) have also been adapted to a multiscale framework (mlcm), which can be used for various applications such as electrolytes and medical simulations (monte_carlo_for_electro; monte_carlo_for_bio). In this paper, we draw inspiration from multiscale methods to introduce a novel and efficient GNN training framework.
@@
In the context of graphs, the recent work in shi2022clustergnn has shown that a matching problem can be solved using a coarse-to-fine strategy, and cai2021graph proposed different coarsening strategies. Other innovations include the recent work of li2024graph, whose goal is to mitigate the lack of labels in semi-supervised and unsupervised settings. We note that it is important to distinguish between algorithms that use a multiscale structure as a part of the network, such as Unet (u_net) and Graph-Unet (graph_u_net), and multiscale training that utilizes multiscale for training, such as the recent work on convolutional neural networks in multiscale_cnn. In particular, the latter, multiscale training, which is the focus of this paper, can be potentially applied to any GNN architecture.
@@ 3 Preliminaries and Notations
[…]
@@
For instance, in the well-known Graph Convolution Network (GCN) (gcn), each layer is defined using Equation (1), such that ${\bf X}^{(l+1)}$ is the output of the $l$-th layer and serves as the input for the $(l+1)$-th layer. For a network with $L$ layers, we denote by ${\boldsymbol{\theta}}=\{{\bf W}^{(1)},\ldots,{\bf W}^{(L)}\}$, the collection of all learnable weight matrices within the network.
@@
Lastly, we note that, as shown in Section˜6, our approach is general and applicable to various GNN architectures.
@@ 4.1 Coarse-to-Fine Training
One seemingly simple idea is to solve the $r$-th surrogate problem in Equation (4) and use the obtained weights to solve the problem on the (r1)(r-1)-th graph, that is ${\cal G}_{{r-1}}$, which is larger. This approach can be beneficial only if $\widehat{\boldsymbol{\theta}}_{r+1}\approx\widehat{\boldsymbol{\theta}}_{r}$, that is, the obtained optimal weights are similar under different scales, and if the total amount of work on the coarse scale is smaller than on the fine one, such that it facilitates computational cost reduction. The multiscale framework is summarized in Algorithm˜1.
@@
While Algorithm˜1 is relatively simple, it requires obtaining the $r$-th surrogate problem by coarsening the graph, as well as a stopping criterion for the $r$-th optimization problem. Generally, in our coarse-to-fine method, we use the following graph coarsening scheme:
@@
Choosing the non-zero entries of the matrix ${\bf P}_{r}$, which are the nodes that remain in the graph after pooling, can be implemented in various ways. For example, ${\bf P}_{r}$ can be obtained by sampling random nodes of the fine graph $\mathcal{G}$, as illustrated in Figure˜1 (left), or with some desired pooling logic, such as choosing nodes of the highest degree in the graph, which is a method often described as $Topk$ pooling (huang2015top) and we present an illustration of it in Figure˜1 (right). The power $p$ in Equation (5) controls the connectivity of the graph, where $p=1$ is the original graph adjacency matrix, and using a value $p>1$ will enhance the connectivity of the graph prior to the coarsening step. We note that the range of applicable choices for $p$ is subject to the given data, as ${\bf A}_{r}$ has to contain fewer edges than the original matrix $\bf A$} in order for the resulting graph to be coarse.
In Algorithm˜1, we apply GNN layers such as in Equation (1), using the coarsened feature matrix ${\bf X}_{r}$ and its corresponding adjacency matrix ${\bf A}_{r}$, that is based on a desired pooling ratio. We note that there is no re-initialization of the weight matrix $\bf W$} when transferring from coarse grids to finer ones, allowing a continuous learning process along the steps of the algorithm. As we show later in our experiments in Section˜6, our coarse-to-fine training approach, summarized in Algorithm˜1, is not restricted to a specific GNN architecture or a pooling technique.
@@ 4.2 Sub-to-Full Training
We now describe an additional method for efficient training on GNNs, that relies on utilizing subgraphs, called Sub-to-Full. At its core, the method is similar to Coarse-to-Fine, as summarized in Algorithm˜1, using the same coarsening in Equation (5). However, its main difference is that we now choose the indices of ${\bf P}_{r}$ such that the nodes ${\bf X}_{r}$ form a subgraph of the original graph $\cal G$.
A subgraph can be defined in a few possible manners. One is by choosing a root node $j$, and generating a subgraph centered around that node. If the graph $\cal{G}$ has some geometric features, meaning each input node corresponds to a point in the Euclidean space $\mathbb{R}^{d}$, one might consider constructing a subgraph with $n_{r}$ nodes to be the $n_{r}$ nodes that are closest to the chosen center node in terms of Euclidean distance norm. Otherwise, we can consider a subgraph created using an ego-network (ego_networks) of khopk-hop steps, such that we consider all neighbors that are $k$ steps away from the central node. An illustration of this method is presented in Figure˜2. The number of hops, $k$, is a hyperparameter of the algorithm process, and we increase it as we go from coarse levels to fine ones within the series of graph $\{\mathcal{G}_{r}\}_{r=1}^{R}$.